The home the Bensons purchased 11 years ago for $100,000 is now appraised at $250,000.

       

What has been the annual rate of appreciation of the value of their home during the 11-year period? (Do not round intermediate calculations and round your final answers to 2 decimal places.)

       
The compound annual rate of appreciation %   

If a $16,135 investment grew to $25,310 in 4.5 years of quarterly compounding, what effective rate of return was the investment earning? (Do not round the intermediate calculations. Round your answer to two decimal places.)

           
Effective rate %

A 180-day promissory note for $2000 bears 7% simple interest. After 60 days it is sold to a bank that charges a simple discount rate of 7%.

a)

Calculate the price paid by the bank for the note. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Price paid by the bank $   
b)

What simple interest rate did the original owner of the note actually earn? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Simple interest rate %
c)

If the note is actually paid on the due date, what simple discount rate will the bank have realized? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Simple discount rate %
Answer #1

(1)

Annual appreciation rate = (250,000 / 100,000)(1/11) - 1 = (2.5)(1/11) - 1 = 1.0869 - 1 = 0.0869 = 8.69%

(2)

Number of quarters = 4.5 x 4 = 18

If nominal quarterly interest rate be r, then

16,135 x (1 + r)18 = 25,310

(1 + r)18 = 1.5686

1 + r = (1.5686)(1/18) = 1.0253

r = 0.0253

Effective annual interest rate = (1 + 0.0253)4 - 1 = (1.0253)4 - 1 = 1.1052 - 1 = 0.1052 = 10.52%

NOTE: As HOMEWORKLIB Answering Policy, 1st 2 questions are answered.